Make 500 ml of 15mM NaCl from a 250 mM NaCl stock


1
Expert's answer
2021-02-25T10:46:22-0500

n (NaCl)1 = n (NaCl)2

CM = n/V

n = CM x V

n (NaCL) = 0.015 x 0.500 = 0.0075 mol

n = m/M

V1 (NaCl) = n/CM = 0.0075/0.25 = 0.03 L = 30 ml


Answer: Take 30 ml of 250 mM solution and make up with distilled water to 500 ml.